OpenStudy (anonymous):

Please help!! Will Fan and Medal! c: 1. Consider x2 + 4x – 5 = y. a) What are the roots of the equation? b) Find the Vertex. c) Write the equation of the axis of symmetry. d) Name the y-intercept.

OpenStudy (anonymous):

y = x^2 + 4x - 5 y = (x + 5) (x - 1) Let y = 0. (x + 5) (x - 1) = 0 x + 5 = 0 x = -5 x - 1 = 0 x = 1 So, (1, -5) will be the answer.

OpenStudy (anonymous):

Hope that helps ^

OpenStudy (anonymous):

thank you c: